Three balls are drawn in succession without replacement from an urn containing 3 red balls, 2 blue balls and 2 yellow balls. Let Z be the random variable representing the number of blue balls.




Sample space (all outcomes)



Possible values



Properties 1 and 2



Probability Mass Function



Histogram

1
Expert's answer
2022-02-18T12:41:18-0500

Solution:

3 red balls, 2 blue balls and 2 yellow balls.

Three balls are drawn in succession without replacement.

Z be the random variable representing the number of blue balls.

(a) S={RRR, RRB, RRY, BBR, BBY, YYR, YYB, BYR}

(b) Possible values of Z = {0,1,2}

(c) Properties 1 and 2 do not make sense. It is incomplete data.

(d) Probability Mass Function:

P(Z=0)=3/8 = 37.5%

P(Z=1)=3/8=37.5%

P(Z=2)=2/8 = 1/4=25%

(e)Histogram: